OpenStudy (cedric):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (10, 35) and (150, 55), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points?

OpenStudy (anonymous):

Remember that slope is the change in y divided by the change in x. \[m = \Delta y / \Delta x\]

OpenStudy (cedric):

ok srry i was eating let me try

OpenStudy (cedric):

slope = (35 - 55)/(10 - 150) = -20/(-140) = 1/7

OpenStudy (cedric):

is that correct?

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

\[M = \frac{y2-y1}{x2-x1} \]

OpenStudy (anonymous):

u have used that formula correctly, so it is correct.

OpenStudy (cedric):

thanks man